# Brill-Noether loci on an Enriques surface covered by a Jacobian Kummer surface

### v1

@@ 3.2. Examples
$(ii)$ For the second part, it is sufficient to prove it only for $D$, since any nonzero effective subdivisor would also be a sum of nodes and the same argument could be applied in the respective case. Write $D=\sum_{i\in I}a_{i}E_{i}$ with $a_{i}>0$ and $E_{i}$ distinct nodes. Take some $j\in I$ and simply note that $(D.E_{j})=-2a_{j}<0$, whereas $(\theta^{*}D.E_{j})\geq 0$.
$(iii)$ Since the nodes don’t intersect, $(D)^{2}=\sum_{i=1}^{n}(E_{i})^{2}=-2n$. ∎
@@
The last proposition remains obviously true if we replace nodes by tropes. Notice that for $(i)$ the same argument works for sums of nodes and tropes, as long as there is no intersection between the nodes and the tropes, or if there exists either a node which intersects no trope, or a trope which intersects no node. For the former case, just choose in the proof of the claim $E_{n}$ to be, for example, that node which intersects no trope.
@@
According to Proposition 3.6, some divisors which satisfy the conditions of Theorem 3.3 should look like $E_{0}+E_{13}+E_{13}$, or $3E_{23}+E_{14}+2E_{56}$. The only thing that we should care about is having at least $3$ terms in our sum (counting multiplicities). The same idea applies for sums of tropes, as mentioned in the previous remark.
For the other types of divisors mentioned in the remark, we need to see how to choose some which fit the other conditions of Theorem 3.3. The last condition is the easiest to handle by far, since we know very well how the intersection product behaves. The second one is a bit trickier, and the following two results will tackle the problem of finding divisors which satisfy it.
@@
Let $D=\sum_{i\in I}a_{i}E_{i}+T$ with the $E_{i}$’s distinct nodes, $a_{i}>0$ and $(E_{i}.T)=0$ for all $i\in I$. Then $D\not\sim\theta^{\ast}D$ if and only if $D$ has one of the following properties:
@@
Let $D=\sum_{i\in I}a_{i}E_{i}+T$ with the $E_{i}$’s distinct, $a_{i}>0$ and $(E_{i}.T)=0$ for all $i\in I$. Then $D$ has no $\theta^{\ast}$-invariant nonzero effective subdivisor if and only if for all $i$, $E_{i}\neq\theta^{\ast}T$.
@@
Let $D$ be as in the corollary with the property that $E_{i}\neq\theta^{\ast}T$ for all $i\in I$. Then, because of the corollary and the previous remark, we know that $D$ satisfies the first two conditions of 3.3. For the last one, notice that $(D)^{2}=\sum_{i\in I}(E_{i})^{2}+(T)^{2}=(-2)\cdot(\#I+1)$, so the third condition is satisfied if and only if #$\#I\geq 2$.

### v2

@@ 3.2. Examples
$(ii)$ For the second part, it is sufficient to prove it only for $D$, since any nonzero effective subdivisor would also be a sum of nodes and the same argument could be applied in the respective case. Write $D=\sum_{i\in I}a_{i}E_{i}$ with $a_{i}>0$ and $E_{i}$ distinct nodes. Take some $j\in I$ and simply note that $(D.E_{j})=-2a_{j}<0$, whereas $(\theta^{*}D.E_{j})\geq 0$ (this is because, as seen in Lemma 2.5 and the table below it, $\theta^{*}D$ will be a sum of tropes and the intersection product between a node and a trope is non-negative).
$(iii)$ This is immediate, since distinct nodes don’t intersect and all the $E_{i}$’s are $-2$ curves. ∎
@@
The last proposition remains obviously true if we replace nodes by tropes. Notice that for $(i)$ the same argument works for sums of nodes and tropes, as long as there is no intersection between the nodes and the tropes.
@@
According to Proposition 3.7, some divisors which satisfy the conditions of Theorem 3.3 could look like $E_{0}+E_{12}+E_{13}$, or $3E_{23}+E_{14}+2E_{56}$. The only thing that we should care about is having at least $2$ terms in our sum (counting multiplicities). The same idea applies for sums of tropes, as mentioned in the previous remark.
Remark 3.10.
A particularly interesting example is produced by the sum of only two nodes, for instance $D=E_{0}+E_{12}$. In this case, by Theorem 3.3 and Remark 3.5, the first Brill-Noether locus $W^{1}_{H}(2;c_{1},c_{2})$ associated to $D$ will be a divisor in the moduli space $M_{H}(2;c_{1},c_{2})$.
For the other types of divisors mentioned in the Remark 3.8, we need to see how to choose some which fit the other conditions of Theorem 3.3. The last condition is the easiest to handle by far, since we know very well how the intersection product behaves. The second one is a bit trickier, and the following two results will tackle the problem of finding divisors which satisfy it.
@@
Let $D=\sum_{i\in I}a_{i}E_{i}+T$ with the $E_{i}$’s distinct nodes, $a_{i}>0$ and $T$ a trope such that $(E_{i}.T)=0$ for all $i\in I$. Then $D\not\sim\theta^{\ast}D$ if and only if $D$ has one of the following properties:
@@
Let $D=\sum_{i\in I}a_{i}E_{i}+T$ with the $E_{i}$’s distinct, $a_{i}>0$ and $T$ a trope such that $(E_{i}.T)=0$ for all $i\in I$. Then $D$ has no $\theta^{\ast}$-invariant nonzero effective subdivisor if and only if for all $i$, $E_{i}\neq\theta^{\ast}T$.
@@
Let $D$ be as in the corollary with the property that $E_{i}\neq\theta^{\ast}T$ for all $i\in I$. Then, because of the Corollary 3.13 and the Remark 3.8, we know that $D$ satisfies the first two conditions of Theorem 3.3. For the last one, notice that $(D^{2})=\sum_{i\in I}a_{i}^{2}(E_{i}^{2})+(T^{2})=(-2)\cdot(\sum_{i\in I}a_{i}^{2}+1)$, so the third condition is also fulfilled.
We gather all that was discussed previously in the following proposition which shows that we have a second family of divisors on the Jacobian Kummer surface satisfying the hypotheses of our main theorem:
Proposition 3.15.
Let $D=\sum_{i\in I}a_{i}E_{i}+T$ with the $E_{i}$’s distinct, $a_{i}>0$ and $T$ a trope such that $(E_{i}.T)=0$ for all $i\in I$. Assume furthermore that for all $i$, $E_{i}\neq\theta^{\ast}T$ Then:
  • (i)
    $h^{0}(\mathcal{O}_{X}(D))=1$;
  • (ii)
  • (iii)
    $(D^{2})\leq-4$.
We finish by mentioning that the divisors described in Propositions 3.7 and 3.15 are not the only ones satisfying the hypothesis of Theorem 3.3, as the following specific example shows:
Example 3.16.
Let us consider the divisor $D=E_{0}+E_{12}+T_{136}+T_{146}$. As we can see from Lemma 2.5, the nodes and the tropes are mutually disjoint. Thus, as mentioned in Remark 3.8, we will have that $h^{0}(\mathcal{O}_{X}(D))=1$. Note that $D$ is not $\theta^{*}$-invariant, because $(E_{0}.D)=-2$, whereas $(E_{0}.\theta^{\ast}D)=(E_{0}.T_{456}+T_{3}+E_{26}+E_{15})=1$ (see Lemma 2.5 and the table describing the action of $\theta^{*}$ on the nodes and the tropes). With Propositions 3.7 and 3.15 in mind, we also see that $D$ has no $\theta^{\ast}$-invariant proper subdivisors. Since $(D^{2})=-16\leq-4$, we conclude that $D$ satisfies the hypothesis of Theorem 3.3.